OpenStudy (anonymous):

hi im doing some math homework and i am having a hard time on this question cansomeone help me? Which of the following equations is equivalent to -4(-2 + w) = 8 ?

OpenStudy (anonymous):

where r the "following"????

OpenStudy (anonymous):

A. 8 = 8 B. 8 − 4w = 8 C. 8 − 4w = -32 D. 8 − 4w = 32

OpenStudy (anonymous):

im sorry i forgot to say the following :)

OpenStudy (anonymous):

B. 8 − 4w = 8 you apply the distributive law on the bracket and open it up.....

OpenStudy (anonymous):

thankyou! :)